OpenStudy (anonymous):

What is the sum of the arithmetic sequence 149, 135, 121, …, if there are 28 terms? help pleasee

OpenStudy (cwrw238):

sum of n terms = (n/2)[2a + (n-1)d] here n = 28, a = 149 (first term), d = common difference common difference = 135 - 149 = -14

OpenStudy (cwrw238):

plug these values in to get your answer

OpenStudy (anonymous):

i got -1120

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

14 * (298 + 27(-14)) = -1120 yes